OpenStudy (anonymous):

Fully simplify the expression: ( 5 / x^2 ) + ( 4 / x^2+x ) Of course x^2 factors as x * x ; be sure to also factor x^2 + x when determining the least common denominator.

OpenStudy (anonymous):

\[\frac{5}{x^{2}} + \frac{4}{x(x+1)} \]

OpenStudy (anonymous):

\[\frac{5x^{2}+5x+4x^{2}}{x^{3}(x+1)}\]

OpenStudy (anonymous):

you finish it, i'm lazy :)

OpenStudy (anonymous):

ps and x on top cancels

OpenStudy (anonymous):

So I got stuck.. after combining the top, I got (9x^2+5x) How do I factor out the bottom? x^3(x+1) :/

OpenStudy (anonymous):

Never mind, got it. x^2(x+1) Thanks!